OpenStudy (anonymous):

a bag contains 4 blue marbles, 5 red, 1 green and 2 black, picking two at random (w/out replacement), what is the probability that the 1st is red and the 2nd is black?

OpenStudy (ryguy):

your name is mean ahahahah

OpenStudy (anonymous):

The probability of picking a red one first is 5/12 as there are 5 red marbles and 12 suming up. Now the possiblity of picking a black one second is 2/11 since there are 2 black and 11 of all marbles since we took out one already. So the probability of both happening is 5/12*2/11=60/132

OpenStudy (anonymous):

so the probability of drawing at least 1 blue is?

OpenStudy (anonymous):

14/33??

OpenStudy (anonymous):

If you pick out two . 4/12+4/11=92/132 Notice how in previous one you multiply and now you add. That is because in previous one you need to have both terms met in two attempts, and in this one only one.

OpenStudy (anonymous):

I see... so the last question is the 2nd is black IF the 1st is green

OpenStudy (anonymous):

Well, now you know there are only eleven marbles. The amount of black one is not changed, only the whole amount of them. So the possibility is 2/11

OpenStudy (anonymous):

Awesome, thanks so much